# Can LLMs Really Learn to Translate a Low-Resource Language from One Grammar Book?

### v1

@@ Appendix A Kalamang Grammar Book Extract
(examples 17 and 18) explaining a morphological feature of Kalamang.
Appendix B Typological Feature Prompt
@@
Appendix C Prompt Examples
@@
Kalamang: =a = English: focus marker Kalamang: a = English: filler Kalamang: a’a = English: yes Kalamang: adat = English: tradition Kalamang: ade = English: pejorative interjection Kalamang: adi = English: interjection of pain Kalamang: adu = English: interjection of surprise or pain Kalamang: afukat = English: avocado Kalamang: ahat = English: sunday |
| Pbook |
@@
Appendix D Prompt Vocabulary Statistics
In Table D, we show test set out-of-vocabulary (OOV) type counts (i.e. unique words) and corresponding test set type coverage in the input prompt for each setting. If the prompt includes a word that is in the test set in the target language, we count that as an in-vocabulary type, and words which do not appear in the prompt as OOV; our denotation of OOV is therefore unrelated to the model’s vocabulary. We additionally include token counts (individual occurrences of types) for each prompt.
@lccccc@
@@ Appendix D Prompt Vocabulary Statistics
Appendix E Test set analysis
To illustrate the weakness of the kgv test set, we generate eng--xxx test sets in: Dutch (nld), German (deu), French (fra), and Spanish (spa) using Google Translate777https://cloud.google.com/translate, and test Gemini’s performance on these sets to find the upper bound. Table 9 shows the test set is weak and a score below 50 ChrF++ falls far below the observed high-resource upper bound. The 100 example set also falls well below standard translation test sets in size, usually 500-1000 examples (Costa-jussà et al., 2024). We addressed the issues of simplicity and size by testing on the Nepali Flores test set.
| Setting | Best | 0-shot | ||||||||||
| Language | kgv | kgv | nld | deu | fra | spa | ||||||
| Direction | ||||||||||||
Appendix F Additional Fine-tuning results
Table F shows translation results for fine-tuning the instruction-tuned Gemini on data, and tested in a 0-shot setting. Included are results with Llama-ft and fine-tuned NLLB. Fine-tuning the small MT model is more effective than tuning an LLM in this particular 0-shot setting.
@lcccccc@
@@ Appendix F Additional Fine-tuning results
Appendix G Statistical Tests
As discussed in Section 5.2, we fit linear regression models to ChrF++ score with test set type coverage as the independent variable for Gemini engkgv translation experiments. We find the models are significantly useful in both directions, (). For eng–kgv: , and for kgv–eng: . The correlations of these results are also significant, where for eng–kgv: , and for kgv–eng: .
Finally, in modelling ChrF++ with prompt tokens as the independent variable for , we find the resulting linear models are not significant. For eng–kgv, ; and for kgv–eng, . There is no correlation between the number of tokens and the observed ChrF++ score.
Appendix H Qualitative Evaluation
@@
Appendix I Limitations
In addition to those noted in the main paper, we acknowledge the following limitations of this work. While we combine the Kalamang test sets to give a 100 example set, this is still far below a standard test set for MT, often 1-2k sentences. In Kalamang, we are limited by the availability of additional data. However we do test Nepali with the Flores devtest set of 1012 examples which provides a more realistic low-resource translation setting. Nepali experiments also address generalisation issues of focusing only on one XLR language. Regarding evaluation, we note that many differences in ChrF++ score were fairly small, and as reported in Kocmi et al. (2024) a difference in ChrF (note, not ChrF++) of 3.05 is required for more than 90% of humans to agree that a system is better than another in practice; this emphasises the need for future experiments and qualitative analyses (see Appendix H for a small scale qualitative analysis).
Further, the majority of our translation experiments are run on Gemini-1.5-Flash, an API-only LLM. Given the nature of our long-context experiments, we are necessarily limited in our choice of model—at the time of running experiments and to our knowledge, no other model family can handle context lengths over 200k tokens which is necessary for the entire Kalamang book. We run selected short-context experiments with the open-weight Llama-3.1-8B model to improve the generalisation of our results, and we leave tests with other long-context models to future work. We finally note that while ideally we would have a larger kgv test set, running long-context inference of paid API models for >1k examples becomes prohibitively expensive. This limitation applies to the entire method of long-context LLM prompting, justifying the fine-tuning of smaller, open-weight, local models for XLR translation instead - especially for members of these language communities who are unlikely to have access to large API models but may have access to free GPUs through services such as Google Colab101010https://colab.research.google.com/ and Kaggle111111https://www.kaggle.com/code.

### v2

@@ Appendix A Kalamang Grammar Book Extract
Appendix B Statistical Tests
As discussed in Section 5.1, we fit linear regression models to ChrF++ score with test set type coverage as the independent variable for Gemini engkgv translation experiments. We find the models are significantly useful in both directions according to the F-test, (). For eng–kgv: , and for kgv–eng: . The Pearson correlations of these results are also significant, where for eng–kgv: , and for kgv–eng: .
Finally, in modelling ChrF++ with prompt tokens as the independent variable for , we find the resulting linear models are not significant according to the F-test. For eng–kgv, ; and for kgv–eng, . There is no correlation between the number of tokens and the observed ChrF++ score.
Appendix C Test set analysis
To illustrate the weakness of the kgv test set, we generate eng--xxx test sets in: Dutch (nld), German (deu), French (fra), and Spanish (spa) using Google Translate555https://cloud.google.com/translate, and test Gemini’s performance on these sets to find the upper bound. Table 6 shows the test set is weak and a score below 50 ChrF++ falls far below the observed high-resource upper bound. The 100 example set also falls well below standard translation test sets in size, usually 500-1000 examples (Costa-jussà et al., 2024). We addressed the issues of simplicity and size by testing the npi and gug Flores test sets.
| Setting | Best | 0-shot | ||||||||||
| Language | kgv | kgv | nld | deu | fra | spa | ||||||
| Direction | ||||||||||||
Appendix D Typological Feature Prompt
@@
Appendix E Prompt Examples
@@
Kalamang: =a = English: focus marker Kalamang: a = English: filler Kalamang: a’a = English: yes Kalamang: adat = English: tradition Kalamang: ade = English: pejorative interjection Kalamang: adi = English: interjection of pain |
| Para |
@@
Appendix F Prompt Vocabulary Statistics
In Table F, we show test set out-of-vocabulary (OOV) type counts (i.e. unique words) and corresponding test set type coverage in the input prompt for each setting. If the prompt includes a word that is in the test set in the target language, we count that as an in-vocabulary type, and words which do not appear in the prompt as OOV; our denotation of OOV is therefore unrelated to the model’s vocabulary. We additionally include token counts (individual occurrences of types) for each prompt.
@lrrrrr@
@@ Appendix F Prompt Vocabulary Statistics
Appendix G Additional Fine-tuning results
Table G shows translation results for fine-tuning the instruction-tuned Gemini on data, and tested in a 0-shot setting. Included are results with Llama-ft and fine-tuned NLLB. Fine-tuning the small MT model is more effective than tuning an LLM in this particular 0-shot setting.
lrrrrrr@
@@ Appendix G Additional Fine-tuning results
Appendix H Limitations
In addition to those noted in the main paper, we acknowledge the following limitations of this work. While we combine the Kalamang test sets to give a 100 example set, this is still far below a standard test set for MT, often 1-2k sentences. In Kalamang, we are limited by the availability of additional data. However we do test Nepali and Guarani with the Flores devtest set of 1012 examples which provides more realistic low-resource translation settings. Nepali and Guarani experiments also address generalisation issues of focusing only on one XLR language. Regarding evaluation, we note that many differences in ChrF++ score were fairly small, and as reported in Kocmi et al. (2024) a difference in ChrF (note, not ChrF++) of 3.05 is required for more than 90% of humans to agree that a system is better than another in practice; this emphasises the need for future experiments and qualitative analyses (see Appendix I for a small scale qualitative analysis).
Further, the majority of our translation experiments are run on Gemini-1.5-Flash, an API-only LLM. Given the nature of our long-context experiments, we are necessarily limited in our choice of model—at the time of running experiments and to our knowledge, no other model family can handle context lengths over 200k tokens which is necessary for the entire Kalamang book. We run selected short-context experiments with the open-weight Llama-3.1-8B model to improve the generalisation of our results, and we leave tests with other long-context models to future work. We finally note that while ideally we would have a larger kgv test set, running long-context inference of paid API models for >1k examples becomes prohibitively expensive. This limitation applies to the entire method of long-context LLM prompting, justifying the fine-tuning of smaller, open-weight, local models for XLR translation instead—especially for members of these language communities who are unlikely to have access to large API models but may have access to free GPUs through services such as Google Colab666https://colab.research.google.com/ and Kaggle777https://www.kaggle.com/code.
Appendix I Qualitative Evaluation